OpenStudy (anonymous):

how do you solve for x in this equation...ln (9-6 x) - 2ln x = ln 3

OpenStudy (aum):

Make use of the following log properties: \(\log(x)^n = n * \log(x)\) \(\log(A) - \log(B) = \log(A/B)\) If \(\log(A) = \log(B)\) then \(A=B\)

OpenStudy (aum):

\[ \ln (9-6 x) - 2\ln x = \ln 3 \\ \ln (9-6 x) - \ln x^2 = \ln 3 \\ \ln((9-6x) / x^2) = \ln(3) \\ (9-6x) / x^2 = 3 \\ 9-6x = 3x^2 \\ 3x^2 + 6x - 9 = 0 \\ x^2 + 2x - 3 = 0 \\ \]Can you solve it now?

OpenStudy (anonymous):

Yeah i got it thanks!

OpenStudy (aum):

You are welcome. Remember you cannot take logarithm of a negative number. If you get an x value that will end up taking log of a negative number discard that solution as extraneous solution.